### v1

@@ 4 Methods
In the following subsections, we describe how we used an agent-based modeling approach to simulate the opinion dynamics of collectives with homophily in space. Then, we provide the metrics we used to evaluate the behavior and performance of the collective.
@@ 4.1 Model
Our baseline model simulates Exploiter agents, whose behavior follows two chief components, conformity and homophily. Thereafter, we introduce a new behavior role for agents, called Messenger. To decide which role to take, agents use DMP, a stochastic Markov process. In the following, we elaborate on our modeling of agents’ behavior.
@@ 4.1.1 Conformity as an opinion-updating rule
Agents are interconnected via the communication network and continuously exchange information with their local neighbors. Conformity, as a form of social influence, causes individuals to change their opinions to get closer to each other when they are exposed to their social neighbors’ opinions. Therefore, conformity poses a constraint on the opinion of agents in the network. Following the DeGroot naive social learning model [11], this dynamic constraint describes how agents modify their opinions based on the information they receive from their neighbors. The updating rule of the opinion is formulated as a weighted average of three different components [41]. These three sources of information influencing each agent’s opinion are individual memory of opinion $(z_{i}^{t}\in\mathbb{R})$, environmental signal $(s_{i}^{t})$, and social signal $(\sum\limits_{j\in\mathbb{N}_{i}}{{z}_{j}^{t}})$ which are described in:
@@ 4.1.2 Homophily as a motion constraint
In our model, agents are not fixed in the information landscape and they actively move in it to search for sources of information matching their opinion. Their movement is determined by the information they receive from the environment and their local neighbors, therefore their neighbors can indirectly make them move in space. This adds to the formation of opinions and drives more complex collective motions in space. To model this, we used homophily as a mechanism for agents to change their position in space so that the information they receive fits better with the average of their local neighbors. This movement is considered an extra step to increase the consensus and happens in the space domain. To implement this movement, we defined an objective function based on the difference of two signals: what comes from the environment, and what the local neighbors agree upon. The difference makes a potential-like function in space that can push agents to specific points where the difference is minimal. So, homophily is an effort to minimize the difference between two values as dissonance value:
$d_{i}^{t}=\frac{1}{2}\left(s_{i}^{t}-z_{\text{loc},i}^{t}\right)^{2}\ .$ (2)
With the local collective average that agent $i$ observes being defined as $z_{\text{loc},i}=\sum_{i=1}^{\mathbb{N}_{i}}z_{i}/N_{i}$. Agents need to optimize this objective function to satisfy the homophily constraint. To implement it in a distributed way, we applied a minimal sample-wise pseudo-gradient descent (same as in [41]), where agents use the differentiation of the samples they measure, as an approximation of the gradient. We used this optimization method since it is independent of the gradient of the objective function, and requires minimal capabilities, being applicable for engineering cases as we showed in [17]. Agents constantly evaluate the dissonance value while they move in space. To approximate the slope of the function at position $\textbf{s}_{i}^{t}=\begin{bmatrix}x_{i}^{t},y_{i}^{t}\end{bmatrix}^{\text{T}}$, agents calculate the differentiation of the objective function over the step they took in the last time step. A decaying memory (weighted by $\beta$) of this differentiation smoothens the approximation of the gradient:
@@
In cases where an agent does not have any neighbors, the movement follows only the random walk. Also, solitary agents will continue walking randomly and update their opinions based on the environmental information until they encounter a neighbor.
4.2 New Behavior: Data Ferrying by Messengers
A potential solution to tackle over-exploitation and the formation of echo chambers due to the limited communication range is to restore the effective connectivity of the network, especially the inter-cluster links of the network. The information can flow across clusters with different opinions and diffuse throughout the network. From an engineering perspective, a trivial solution would be to scale the problem by increasing the communication range, hence pushing the system to higher connectivity regimes. Improving the hardware, if possible, comes with physical constraints and increases the cost of the designed system. However, by harnessing the mobility of the agents, an alternative solution is to transmit information via the physical movement of agents carrying it in space. This way, different clusters can exchange information over distances that are possibly much larger than the communication range. To achieve this, we introduce a new, so-called “Messenger” state for agents to transfer the information as they move in space. A Messenger can be seen as embodied data that moves around in space and shares the information with its local neighbors that it encounters on its way. This is a similar concept as Zealots or stubborn agents who do not change their opinions [43, 44, 45]. A Messenger moves in space and establishes new links with other agents. The Messenger state should have the following fundamental properties:
@@ 4.2 New Behavior: Data Ferrying by Messengers
In other words, a Messenger agent establishes long-distance uni-directional links by broadcasting its opinion to local neighbors while moving randomly in space. A Messenger migrating by chance from one cluster to another resembles and implements long, weak ties in a dynamic spatial network [46]. A Messenger state is in contrast with the “Exploiter” state. An Exploiter integrates the information following the model we explained in the previous subsections, whereas a Messenger behaves as a moving memory of information. A receiver agent, whether Messenger or Exploiter, does not distinguish the transmitter of the incoming information. An Exploiter receives social information from other Exploiters or Messengers and integrates it regardless of its source. In contrast, Messengers do not process the incoming information and hence are regarded as information; information about the history of opinions. In this regard, a Messenger provides a link to the history of its opinion or, in some cases, the opinion of its cluster.
Agents can switch back and forth between the two states. When an Exploiter turns into a Messenger, it carries its latest opinion. If this opinion was the local consensus, then the data represents the opinion of the cluster. Similarly, a Messenger can switch back to an Exploiter state. In that case, the agent forgets its data and replaces it with the current information it receives from the environment. This causes another functional benefit of Messengers which is increased exploration. The challenge is to define a mechanism for agents to switch their states. In the following subsection, we propose a simple switching mechanism in a fully decentralized manner based on the DMP. First, we briefly review the properties of DMPs.
4.3 Dichotomous Markov Process (DMP)
@@
This specific type of Markov process provides a probabilistic mechanism for switching between two contrasting states, Exploiter and Messenger in our case. The ability to implement it in a decentralized way is an essential feature in our scenario. We implement the DMP such that each agent $i$ at time step $t$ either remains in its current state or switches to the other state. An agent in an Exploiter state switches to Messenger with a transition probability $p_{\text{M}}$, and vice versa with probability $p_{\text{E}}$, per time step. The complement of each transition probability ($1-p_{\text{M}}$, and $1-p_{\text{E}}$) determines the probability of remaining in the current state. The process is illustrated as a state machine in Fig. 9. We implement the process in a discrete-time domain, and we use $\Delta t=1$. Hence, we formulate the DMP in terms of transition probabilities instead of transition rates [42]. The properties of the behavior for each agent are defined by the two parameters  $p_{\text{M}}$,  $p_{\text{E}}$. Two properties that are of interest in this paper are the sojourn time and the expected ratio of Messengers. Assuming the stationary process, the sojourn time [47], denoted as the time between two consecutive switches, for each of the states is defined as follows:
@@
Then, we define sojourn time ($\tau_{\text{S}}$), the average time steps spent in either of the states before switching, as the average of the two sojourn times:
@@ 4.4 Metrics and Setup
Here we define evaluation metrics to benchmark the collective performance. We have metrics of two types: the first type evaluates the collective in the opinion domain, i.e. by using the internal opinion of each agent; while the second does not require access to the internal state of the agents, but their position in the physical space. By quantifying the precision of the collective opinion, the spatial arrangement of agents in space, and the connectivity of the communication network, these metrics capture all relevant aspects of the collective performance. To show how Messengers change the performance of the collective, we normalize the absolute metrics to that of the baseline setup, where there is no Messenger in the collective. Please note that in our finite size stochastic simulations, the expected ratio of Messengers for the bottom right corner of the parameter space is $\mathbb{E}[m]<10^{-8}$, which validates the zero-Messenger assumption. We refer to these metrics as normalized metrics.
@@ 4.4.1 Precision Error
To show the precision of the spatial positioning of the collective, i.e. spatial consensus, we use the same metric and take the position of agents instead. This metric directly quantifies the performance in terms of contour capturing [17]. The precision error in the space domain measures how close are agents to the collective average with regard to the information distribution. We distinguish the two domains by a superscript and define the precision error in the space domain as:
@@ 4.4.2 Number of Clusters
Another metric that measures the cohesion of the agents in the physical space, particularly considering their communication network, is the number of connected components they form in their network. To count the clusters, we find connected components based on the communication range of agents and assign adjacent agents to the same cluster. We count how many clusters are present in the network.
@@ 4.5 Simulation Configuration
As an extension of our previous work [41], we used the same parameters, except for the communication range. To put the system in a regime where echo chambers can potentially emerge, we set the communication range to $r_{\text{comm}}=0.1$. This is similar to the configuration of Kilobots [48], a robotic platform used to study collective behaviors, which we used to implement the baseline setup in a real-world setting [17]. Also, we distributed the probabilities ($p_{\text{E}}$, $p_{\text{M}}$) exponentially, same as in Fig. 4. This way, we can reveal a wide spectrum of the DMP dynamics on various scales. The results that we report are the average of 24 independent Monte-Carlo simulations. Otherwise noted, we used the parameters of the model as reported in Table 1.
@@
$\beta$ Decaying Factor for Gradient Descent 0.9
@@ 5 Data Availability
The datasets generated and analyzed during the current study will be available online.
@@ 6 Code Availability
The underlying code for this study is available in the GitHub repository and can be accessed via this link https://github.com/mohsen-raoufi/messengers.
@@ Acknowledgments
This study was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. The funder played no role in study design, data collection, analysis and interpretation of data, or the writing of this manuscript.

### v2

@@ 4 Methods
In the following subsections, we provide details, first on our agent-based modeling approach to simulate opinion dynamics of collectives with homophilic interactions in space. Then, we introduce the metrics we used to evaluate the behavior and performance of the collective. Finally, we provide information on the parameters used in our simulation.
@@ 4.1 Model
We present first the components of the baseline model that simulates Exploiter agents, whose behavior is determined by two factors: conformity and homophily (see also [1]). Then, we introduce a new behavior role for agents, which we refer to as Messenger, with the switching of agents between the two states implemented via a DMP.
@@ 4.1.1 Conformity as an opinion-updating rule
Agents are interconnected via the communication network and continuously exchange information with their local neighbors. Conformity, as a form of social influence, causes individuals to adapt their opinions to reduce their disagreement with others, when they are exposed to their social neighbors’ opinions. Therefore, conformity poses a constraint on the opinion of agents in the network. Following the DeGroot social learning model [12], this dynamic constraint describes how agents modify their opinions based on the information they receive from their neighbors. The updating rule of the opinion is formulated as a weighted average of three different components [39]: private memory of opinion $(z_{i}^{t}\in\mathbb{R})$, environmental signal $(s_{i}^{t})$, and social signal $(\sum\limits_{j\in\mathbb{N}_{i}}{{z}_{j}^{t}})$ which are described in:
@@ 4.1.2 Homophily as a motion constraint
In our model, agents are not fixed in the information landscape but actively move in it to search for sources of information matching their opinions. Similarly, homophily can be seen as the effort to move and find neighbors that match the information the agent receives. This movement is determined by the information agents receive from the environment and their local neighbors, therefore their neighbors can indirectly induce their movement in space. This adds to the formation of opinions and drives more complex collective motions in space. To model this, we used homophily as a mechanism for agents to change their position in space so that the information they receive fits better with the average of their local neighbors. This movement is considered an extra step to increase consensus and is performed in the spatial domain. To implement this movement, we defined an objective function based on the difference of two signals: what the input from the environment is, and what the local neighbors agree upon. The difference generates a potential-like function in space that biases the agents to move to specific points where the difference is minimal. So, homophily is an effort to minimize the dissonance as the difference between two values:
$d_{i}^{t}=\frac{1}{2}\left(s_{i}^{t}-z_{\text{loc},i}^{t}\right)^{2}\ .\ ,$ (2)
with $z_{\text{loc},i}=\sum_{i=1}^{\mathbb{N}_{i}}z_{i}/N_{i}$ being the local collective average that agent $i$ observes. Agents need to optimize this objective function to satisfy the homophily constraint. To implement it in a distributed way, we applied a minimal sample-wise pseudo-gradient descent (same as in [39]), where agents use the differentiation of the samples they measure, as an approximation of the gradient. We used this optimization method since it is independent of the gradient of the objective function, and requires minimal capabilities, being applicable for engineering cases as we showed in [1]. Agents constantly evaluate their dissonance value while they move in space. To approximate the slope of the function at position $\textbf{s}_{i}^{t}=\begin{bmatrix}x_{i}^{t},y_{i}^{t}\end{bmatrix}^{\text{T}}$, agents calculate the difference of the objective function over the step they took in the last time step. A decaying memory (weighted by $\beta$) of this differentiation smoothens the approximation of the gradient:
@@
In cases where an agent does not have any neighbors, the movement follows only the random walk. Solitary agents will continue walking randomly and update their opinions based on the environmental information until they encounter a neighbor.
4.1.3 Model extension: Data Ferrying by Messengers
A potential solution to tackle over-exploitation and the formation of echo chambers due to the limited communication range is to restore the effective connectivity of the network, especially the inter-cluster links of the network. The information can flow across clusters with different opinions and diffuse throughout the network. From an engineering perspective, a trivial solution would be to scale the problem by increasing the communication range, hence pushing the system to higher connectivity regimes. Improving the hardware, if possible, comes with physical constraints and increases the cost of the designed system. However, by harnessing the mobility of the agents, an alternative solution is to transmit information via the physical movement of agents carrying it in space. This way, different clusters can exchange information over distances that are possibly much larger than the communication range. To achieve this, we introduce a new, so-called ‘Messenger’ state for agents to transfer the information as they move in space. A Messenger can be seen as embodied data that moves around in space and shares the information with its local neighbors that it encounters on its way. This is a similar concept as Zealots or stubborn agents who do not change their opinions [47, 48, 49]. A Messenger moves in space and establishes new links with other agents. The Messenger state should have the following fundamental properties:
@@ 4.1.3 Model extension: Data Ferrying by Messengers
In other words, a Messenger agent establishes long-distance uni-directional links by broadcasting its opinion to local neighbors while moving randomly in space. A Messenger migrating by chance from one cluster to another resembles and implements long, weak ties in a dynamic spatial network [50]. The Messenger state contrasts the “Exploiter” state. An Exploiter integrates the information following the model we explained in the previous subsections, whereas a Messenger behaves as a moving memory of information. A receiving agent, whether Messenger or Exploiter, does not distinguish the transmitter of the incoming information. An Exploiter receives social information from other Exploiters or Messengers and integrates it regardless of its source. In contrast, Messengers do not process the incoming information and can be regarded as mobile carriers of information; the information about the history of opinions. In this regard, a Messenger provides a link to the history of its opinion or, in some cases, the opinion of its cluster.
Agents can switch back and forth between the two states. When an Exploiter turns into a Messenger, it carries its latest opinion. If this opinion corresponds to the local consensus, then the data represents the opinion of the cluster. Similarly, a Messenger can switch back to an Exploiter state. In that case, the agent forgets its data and replaces it with the current information it receives from the environment. This causes another functional benefit of Messengers which is increased exploration.
4.1.4 Switching between states: Dichotomous Markov Process (DMP)
@@
This specific type of Markov process provides a probabilistic mechanism for switching between two contrasting states, Exploiter and Messenger in our case. It can be easily implemented in a decentralized way and requires only minimal computational capabilities per agent. We implement the DMP such that each agent $i$ at time step $t$ either remains in its current state or switches to the other state. An agent in an Exploiter state switches to Messenger with a transition probability $p_{\text{M}}$, and vice versa with probability $p_{\text{E}}$, per time step. The complement of each transition probability ($1-p_{\text{M}}$, and $1-p_{\text{E}}$) determines the probability of remaining in the current state. The process is illustrated as a state machine in Fig. 9. Note, in general, the DMP is formulated in continuous time in terms of transition rates. The probability of switching per time step is then simply the product of the corresponding transition rate and the time step. We implement the process in a discrete-time domain with $\Delta t=1$. Hence, we formulate the DMP directly in terms of transition probabilities instead of transition rates [46]. The properties of the behavior for each agent are defined by the two parameters  $p_{\text{M}}$,  $p_{\text{E}}$. Two properties that are of interest in this paper are the sojourn time and the expected ratio of Messengers. Assuming the stationary process, the sojourn time [51], denoted as the time between two consecutive switches, for each of the states is defined as follows:
@@
Then, we define average sojourn time ($\tau_{\text{S}}$), as the average time spent in either of the states before switching, as the average of the two sojourn times:
@@ 4.2 Metrics and Setup
Here we define evaluation metrics to benchmark the collective performance. We have metrics of two types: the first type evaluates the collective in the opinion domain, i.e., by using the internal opinion of each agent; while the second does not require access to the internal state of the agents, but their position in the physical space. By quantifying the precision of the collective opinion, the spatial arrangement of agents in space, and the connectivity of the communication network, these metrics capture all relevant aspects of the collective performance. To show how Messengers change the performance of the collective, we normalize the absolute metrics to that of the baseline setup, where there is no Messenger in the collective. Please note that in our finite-size stochastic simulations, the expected ratio of Messengers for the bottom right corner of the parameter space is $\mathbb{E}[m]<10^{-8}$, which validates the zero-Messenger assumption. We refer to these metrics as normalized metrics.
@@ 4.2.1 Precision Error
To show the precision of the spatial positioning of the collective, i.e., spatial consensus, we use the same metric and take the position of agents instead. This metric directly quantifies the performance in terms of contour capturing [1]. The precision error in the space domain measures how close are agents to the collective average with regard to the information distribution. We distinguish the two domains by a superscript and define the precision error in the space domain as:
@@ 4.2.2 Number of Clusters
Another metric that measures the cohesion of the agents in the physical space, particularly considering their communication network, is the number of disconnected clusters. To quantify the number of clusters, we identify and count the connected components of the spatial communication network, determined by the communication range of agents.
@@ 4.3 Simulation Configuration
As an extension of our previous work [39], we used the same parameters, except for the communication range. To put the system in a regime where echo chambers can potentially emerge, we set the communication range to $r_{\text{comm}}=0.15$. This is similar to the configuration of Kilobots [52], a robotic platform used to study collective behaviors, which we used to implement the baseline setup in a real-world setting [1]. Also, we distributed the probabilities ($p_{\text{E}}$, $p_{\text{M}}$) exponentially, same as in Fig. 4. This way, we can reveal a wide spectrum of the DMP dynamics on various scales. The results that we report are the average of 24 independent Monte-Carlo simulations. Otherwise noted, we used the parameters of the model as reported in Table 1.
@@
$\beta$ Decaying Factor for Gradient Descent 0.5
@@ 5 Data Availability
The datasets generated and analyzed during the current study are available in this repository: https://depositonce.tu-berlin.de/handle/11303/22195.
@@ 6 Code Availability
The underlying code for this study is available in the GitHub repository and can be accessed via this link: https://github.com/mohsen-raoufi/messengers.
@@ Acknowledgments
This study was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC 2002/1 “Science of Intelligence” – project number 390523135. The funder played no role in the study design, data collection, analysis and interpretation of data, or the writing of this manuscript.
Author contributions
M.R. contributed to conceiving and implementing models and experiments, as well as to analyzing the results. H.H. and P.R. equally supervised the project. M.R. wrote the first draft of the manuscript with input from all authors. All authors reviewed and revised the manuscript.
Competing Interests
All authors declare no financial or non-financial competing interests.
@@
Temporal Development of Transient DMP
As discussed in Sec. 2.4.1, having a non-stationary initial condition of the DMP changes the locally optimal configurations depending on the final time of the simulation. This is shown in Fig. 11 for various snapshots of the opinion precision error at different times.
(a) t=2k
(b) t=4k
(c) t=10k
(d) t=20k
(e) t=40k
Figure 11: Time development of parameter space for precision error of a transient DMP. Here we start the simulation with a collective of all Exploiters. The locally optimal regions shift toward slower DMP configurations over time.
One-dimensional Information landscape